OpenStudy (anonymous):

which of the following is an x-intercept of the function, f(x)= x^3-5x^2-8x+48 A.4 B.-2 C.-4 D.2

OpenStudy (welshfella):

at an x intercept f(x) = 0 so plug in the options into the expression for f(x) and see which one works out to be 0

OpenStudy (welshfella):

for example for x = 4 we get (4)^3 - 5(4)^2 - 8(4) + 48 = ?

OpenStudy (welshfella):

64 - 80 -32 + 48 = ?

OpenStudy (welshfella):

does that come to 0? if it does you have your answer

OpenStudy (anonymous):

I did all four and none of them equal zero unless I did it wrong

OpenStudy (anonymous):

92 44 -28 52

OpenStudy (anonymous):

those were my answers...

OpenStudy (welshfella):

64 - 80 - 32 + 48 = 64 + 48 - 80 - 32 = 112 - 112 = 0

OpenStudy (welshfella):

A is the correct option

OpenStudy (anonymous):

ok thank you. I couldn't get it

OpenStudy (welshfella):

yw

OpenStudy (welshfella):

with something like this it might be easier to do all additions first then subtractions last